OpenStudy (anonymous):

1.Complete the table below to solve the equation 2.5x − 10.5 = 64(0.5x). x f(x) = 2.5x − 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (anonymous):

@imqwerty @Chumpian @MasonP @Lollygirl217 @doulikepiecauseidont @k142 @pooja195 @AnimalLover8 @Empty @JadedInsomniac @jasmine_15 @jamesr @lυἶცἶ0210

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@imqwerty @pooja195 @Empty

imqwerty (imqwerty):

f(2) = 2.5*2 - 10.5 = -5.5 g(2) = 64(0.5*2) = 64 f(3) = 2.5*3 - 10.5 = -3 g(3) = 64(0.5*3) = 96 can u do the rest urself?